# CroCoDai: A Stablecoin for Cross-Chain Commerce

### v1

@@ 2.2 The Dai Stablecoin System (DSS)
In this section, we describe the core features of the existing single-chain DSS. Although there are documents and code repositories that describe the DSS in its various stages of evolution, e.g., [38, 39, 40, 44, 3, 10], we focus on the GitHub code [40] and the description of the multi-collateral DSS [39].
@@
Dai Tokens. In the DSS, Dai tokens are minted by users who deposit cryptocurrency tokens as collateral. In the first iteration of the Dai stablecoin (the single-collateral Dai or “Sai”), only ETH tokens could be deposited for this purpose. The procedure for a user, Alice, who wants to mint new Sai tokens is as follows. First, Alice must send a transaction to a Dai vault to create a Collateralized Debt Position (CDP) [38]. Next, Alice must send ETH tokens as collateral to the contract to fund the CDP. If the CDP is funded, then Alice can use it to create Dai stablecoins and withdraw them from the CDP to her wallet, incurring debt. At any later time, she can send Dai tokens back to the CDP to repay her debt. The above can be generalized to a multi-collateral (but single-chain) DSS by associating each CDP with a token type. Debt increases exponentially over time, and the rate at which it increases is a system parameter called the stability fee.
@@
Liquidations. At all times, Dai CDPs must be overcollateralized, i.e., the ratio of the total value of the Alice’s collateral to her debt must exceed some threshold called the liquidation ratio [3, 38]. By default, the liquidation ratio in Dai is set to $1.5$ – i.e., Alice can withdraw Dai stablecoins worth at most $\frac{2}{3}$ of the total value of the collateral in their CDP [44]. Similarly, Alice can return Dai and withdraw collateral from her CDP as long as the above condition holds. If a CDP has no associated debt, then it can be closed and all remaining collateral can be withdrawn. If, at any time, a CDP’s ratio of its collateral value to its debt is insufficient, then Alice’s CDP is underwater and can be liquidated. CDPs can go underwater either because the value of the collateral has dropped due to price fluctuations, or because the stability fee has caused the debt to become too high. To avoid liquidation, Alice must either deposit additional collateral, or raise Dai to clear the position when it is at risk – if this fails, a collateral auction can be started by other users. During a collateral auction, other users can bid to buy (portions of) Alice’s collateral using Dai stablecoins. Once enough Dai has been raised to repay the CDP’s total debt, the remaining collateral (if any) is sent back to Alice, minus a penalty.
@@
Price Feeds. To determine the value of deposited collateral, MakerDAO uses information from price feeds supplied by oracles [35]. The number of price feeds used by the DSS has increased from 14 between 2017 and November 2019 [26] to 20 currently [43]. Smart contracts that maintain CDPs – vaults – periodically update token prices by setting each price equal to the median of a set of recent prices for the associated token.
Dai Savings Rate. Due to the stability fee, the total amount of debt in the protocol may grow to exceed the total number of stablecoins in circulation. For example, if Alice’s debt of 100 Dai has grown to 105 Dai, then clearing her debt and closing the CDP would (in principle) leave 5 Dai in the vault. These surplus Dai are part of the Dai Savings Rate (DSR) mechanism, which allows users who deposit stablecoins in a DSR contract to periodically withdraw surplus Dai tokens from associated vaults. The DSR helps MakerDAO to keep the value of Dai soft-pegged to the dollar by reducing or decreasing the interest rate: if Dai becomes less valuable than the dollar, then the interest rate is increased to incentivize users to buy Dai to deposit in the DSR contract (and decreased for higher Dai values).
@@
Governance. All major decisions about the functioning of the protocol are taken by users who hold MKR, which is MakerDAO’s governance token. Such decisions include 1) verifying and adding smart contracts that implement any of the functionalities above, 2) changing the Dai savings rate, and 3) occasionally changing other system-level parameters. The following are the main system-level parameters (see page 15 of [39]): 1. the debt ceiling for each token type, i.e., the maximum amount of total debt that can be created using a given collateral type, 2. the stability fee, 3. the liquidation ratio, 4. the liquidation penalty, 5. the duration of collateral auctions, 6. the duration of auction bids, and 7. the minimum auction step size (which prevents bids that are only marginally above the previous bid, which would waste transaction fees).
@@
Out-of-Scope Concepts. Some of the concepts in the DSS are beyond the scope of this paper. In particular, these are 1) debt auctions, which occur when a collateral auction fails to raise enough Dai to cover the CDP’s debt, 2) surplus auctions, which occur when the amount of Dai left in vaults due to auctions and stability fees exceeds some threshold, and 3) (emergency) shutdowns, which can be triggered after a catastrophic event (e.g., system failure) and during which the total remaining collateral is distributed to Dai holders.
2.3 Existing Cross-Chain Approaches
Cross-chain deals [27] provide a modeling framework to describe sequences of cross-chain token transfers. A cross-chain deal consists of five phases: clearing, escrow, transfer, validation, and commitment. The goal of cross-chain deals is to guarantee atomicity, i.e., that either all transfers occur or none do. In [27], two frameworks for cross-chain deals are proven to satisfy technical safety (all honest users end up with an acceptable payoff) and liveness (no assets are escrowed forever, and if all users are honest then all transfers eventually happen): 1. hashed time lock contracts, 2. dedicated blockchains for commit messages . In cross-chain deals, the cost of a participant abandoning the protocol before completion (sore-loser attacks [56]) can be made explicit – this form of misbehavior can be disincentivized through a premium [18] paid by the parties that abandon the protocol before termination.
Cosmos [7] and Polkadot [33] are blockchain platforms that were specifically developed with blockchain interoperability in mind, but which do not trivially interoperate with existing blockchains outside their ecosystems. Furthermore, Xclaim [58] and zkBridge [54] claim to provide secure cross-chain transfers, but have limitations when transferring to/from smaller blockchain platforms as discussed in the introduction.
3 System Model
3.1 Entities
Our system model has the following types of entities. In the following, we use the notation $[n]$ for $\{1,\ldots,n\}$. A list of symbols can be found in Table 4 in Appendix A.
  • Token: a type of cryptocurrency or an NFT (although typically the former due to the better availability of price information for cryptocurrencies). We denote the number of different tokens in the system by $M\geq 2$.
  • Governance Token: this is a type of cryptocurrency that gives holders the right to vote on decisions that affect the functioning of the crypto-backed stablecoin. The governance token cannot be used as collateral for the stablecoin.
  • Blockchain (or chain): an immutable, distributed ledger that tracks the ownership of its tokens. We assume that each chain either supports smart contracts or that a wrapped version of its native tokens exist on a chain that does (e.g., WBTC [53]). We denote the number of supported chains by $B\geq 2$, and the chain of token $m\in[M]$ by $b_{m}\in[B]$.
  • Consensus Node: is responsible for maintaining the blockchain, e.g., through the creation and/or validation of new blocks. The nature of the consensus nodes depends on the blockchain’s consensus protocol – examples include variations of proof-of-work, proof-of-stake, or Byzantine Fault Tolerance (BFT) protocols. Each consensus node has some amount of consensus power – e.g., processing power in proof-of-work, staked tokens in proof-of-stake, or a private key in a BFT protocol.
  • User: owns blockchain tokens and seeks to exchange them for tokens on another chain. We denote the number of users in our system by $U\geq 1$.
  • Smart contract: a software program on a blockchain that may hold tokens which can be sent to other contracts or wallets depending on the program’s internal logic.
  • Vault: a smart contract or address that holds tokens as collateral, and to/from which tokens can be sent/received leading to the possible creation/destruction of stablecoins managed by another smart contract.
  • Price Oracle: a type of user who provides price information to the system [35], e.g., Chainlink [5]. We denote the number of price oracles in our system by $O\geq 1$. Each oracle $o\in[O]$ can provide price information for any selection of tokens. Let $p_{m\,o\,t}$ be the price of token $m\in[M]$ reported by oracle $o\in[O]$ at time $t\geq 0$. A price feed is a sequence of prices for a given token $m$ reported by an oracle $o$.
3.2 Token Price Model
[…]
To model the evolution of token prices over time, we use Geometric Brownian Motion (GBM). This is a common choice in the financial literature on asset prices [47] and related work on stablecoins, e.g., by Salehi et al. [48], particularly due to the GBM’s role in the Black-Scholes formula for option prices. A stochastic process $S_{t}$ that follows GBM is defined using a stochastic differential equation:
$dS_{t}=\mu S_{t}dt+\sigma S_{t}dW_{t},$ (1)
where $\mu$ determines the drift of $S_{t}$ (i.e., its expected change per unit of time), $\sigma$ represents the volatility of $S_{t}$ (i.e., the variance of the change), and $W_{t}$ is the Wiener process that represents standard Brownian motion [47]. In this model, each token price process $p_{m\,t}$, $t\geq 0$, can be represented using a process defined as in Equation 1. Essentially, this means that the logarithm of the relative difference of token prices, $\log\left(p_{m\,t+t^{\prime}}/p_{m\,t}\right)$, is normally distributed for all $t^{\prime}>0$ with mean and variance dependent on $t^{\prime}$. Unlike [48], we consider a setting with multiple assets; i.e., we have a stochastic process $S_{m\,t}$ for each token $m$ with its own mean $\mu_{m}$ and variance $\sigma^{2}_{m}$. In this setting, we must also consider the covariance between the token prices [47], as cryptocurrency prices are often highly correlated – we will discuss this in more detail in Section 5. More details about how the standard GBM can be used to simulate asset prices can be found in Appendix B.
The standard GBM as decribed above is based on the normal distribution, i.e., $\log(S_{t+t^{\prime}}/S_{t})$ has a normal distribution. However, it is known from the academic literature that the probability distribution of stock price fluctuations is usually more heavy-tailed and that Student’s $t$-distribution is often a better fit [29, 34]. This is also visible in the data for price changes in our dataset – see Figure 5 in Appendix C. As such, we will focus on the $t$-distribution in our analysis. Although the $t$-distribution does not allow for an analytical solution for the asset price change after each individual time slot (e.g., as per (5) in Appendix B), we can use Monte Carlo simulation to analyze its behavior, as we discuss in more detail in Section 5.
3.3 DSS Execution Model
Let $\mathcal{J}_{t}$ denote the set of all CDPs at time $t$, and $c_{j\,t}$ the total number of collateral tokens held in CDP $j\in\mathcal{J}_{t}$ at time $t$. Let $D_{j\,t}$ be the outstanding debt resulting from Dai withdrawals/deposits from/to CDP $j\in\mathcal{J}_{t}$. Each CDP is associated with a specific token type – let $y_{j}$ be the token type of CDP $j$. Debt is subject to fees, i.e., $D_{j\,t+1}=(D_{j\,t})^{1+\delta}$, where $\delta$ represents the stability fee. We denote the number of stablecoins held by user $u\in[U]$ at time $t$ by $h_{u\,t}$. We note that, in principle, $\sum_{u\in[U]}h_{u}\leq\sum_{j\in\mathcal{J}}D_{j\,t}$, as stablecoins are created and transferred to a user account either through withdrawal from a CDP (in which case the created stablecoins and debt are equal) or from a vault through the Dai savings rate (in which case the created stablecoins cannot exceed the excess debt). However, we will consider attacks in Section 3.4 that can cause this assumption to be violated.
We formalize the liquidation criterion as follows: for each CDP $j\in\mathcal{J}_{t}$ and $t\geq 0$, it must hold that $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma,$ where $\gamma$ is the liquidation ratio – if this condition does not hold, then CDP $j$ can be liquidated through a collateral auction. We assume that under standard conditions (i.e., the absence of an attack or sudden price collapse), collateral auctions are concluded successfully, i.e., a sufficient amount of stablecoins is raised to close the position. No user can create debt if the result would violate the liquidation criterion. We also note that price fluctuations do not affect the liquidation criterion, nor does the Dai savings rate as the proportional effect on all CDP debts is the same.
We denote the debt ceiling for each token type $m\in[M]$ by a fraction $\zeta_{m}\in[0,1]$, i.e., the total debt in CDPs of type $m$ relative to the system’s total debt cannot exceed $\zeta_{m}$. (Debt ceilings are specified in absolute terms in DSS contracts, but we assume that the difference is negligible for short periods, e.g., for escrow.) The debt ceilings are used to enforce $\sum_{\begin{subarray}{c}j\in\mathcal{J}_{t}\\ y_{j}=m\end{subarray}}D_{j\,t}<\zeta_{m}\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$ for all $m\in[M]$. Let $D_{t}^{*}=\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$ and $C^{*}_{t}=\sum_{j\in\mathcal{J}_{t}}p_{y_{j}\,t}c_{j\,t}$ denote the total debt and the total collateral value across the system at time $t$, respectively.
3.4 Threat Model
We make the following assumptions about the system entities. We assume that a majority of the governance token holders are honest. For each underlying blockchain, we assume that it is either compromised or not. If it is not compromised, we assume that it guarantees safety and liveness for all of its transactions. However, both safety and liveness can be violated in a compromised blockchain – in particular, transactions in which collateral is sent to/from a CDP can be overturned. For the price oracles, we assume that a majority is online and honest at all times, but that dishonest oracles can report arbitrary prices.
3.5 System Requirements

### v4

@@ 2.2. The Dai Stablecoin System (DSS)
In this section, we describe the existing single-chain DSS. Although there are documents and code repositories that describe the DSS in its various stages of evolution, e.g., (daiwhitepaper, ; multicollateraldai, ; daigithub, ; moin2020sok, ; bhat2021simulating, ; chen2022understanding, ), we focus on the GitHub code (daigithub, ) and the description of the multi-collateral DSS (multicollateraldai, ).
@@
Dai Tokens. In the DSS, Dai tokens are minted by users who deposit cryptocurrency tokens as collateral. In the first iteration of the Dai stablecoin (the single-collateral Dai or “Sai”), only ETH tokens could be deposited for this purpose. The procedure for a user, Alice, who wants to mint new Sai tokens is as follows. First, she must send a transaction to a Dai vault to create a Collateralized Debt Position (CDP) (daiwhitepaper, ). Next, she must send ETH tokens as collateral to the contract to fund the CDP. Alice can use a funded CDP to create Dai stablecoins and withdraw them to her wallet, incurring debt. At any later time, she can send Dai tokens back to the CDP to repay her debt. The above can be generalized to a multi-collateral (but single-chain) DSS by associating each CDP with a token type: in the DSS, this could theoretically include any ERC20 token, including wrapped tokens such as WBTC. Debt increases exponentially over time, and the rate at which it increases is determined by the stability fee.
@@
Liquidations. At all times, Dai CDPs must be overcollateralized, i.e., the ratio of the total value of the Alice’s collateral to her debt must exceed some threshold called the liquidation ratio (bhat2021simulating, ; daiwhitepaper, ). By default, the liquidation ratio in Dai is set to $1.5$ – i.e., Alice can withdraw Dai stablecoins worth at most $\frac{2}{3}$ of the total value of the collateral in their CDP (moin2020sok, ). Similarly, Alice can return Dai and withdraw collateral from her CDP as long as the above condition holds. If a CDP has no associated debt, then it can be closed and all remaining collateral can be withdrawn. If, at any time, a CDP’s ratio of its collateral value to its debt is insufficient, then Alice’s CDP is underwater and can be liquidated. CDPs can go underwater either because the value of the collateral has dropped due to price fluctuations, or because the stability fee has caused the debt to become too high. To avoid liquidation, Alice must either deposit additional collateral, or raise Dai to clear an at-risk position – if this fails, a collateral auction can be started by other users. During a collateral auction, other users can bid to buy (portions of) Alice’s collateral using Dai tokens. Once the CDP’s debt has been cleared, Alice receives the remaining collateral (if any) minus a penalty.
@@
Price Feeds. To determine the value of deposited collateral, MakerDAO uses information from price feeds supplied by oracles – e.g., Chainlink or Uniswap contracts (liu2021first, ). The number of price feeds used by the DSS has increased from 14 between 2017 and November 2019 (gu2020empirical, ) to 20 in late 2023 (mips, ). Smart contracts that maintain CDPs – i.e., the vaults – periodically update token prices by setting each price equal to the median of a set of recent prices for the associated token.
@@
Governance. All major decisions about the functioning of the protocol are taken by users who hold MKR, which is MakerDAO’s governance token. Such decisions include 1) verifying and adding smart contracts that implement any of the functionalities above, 2) changing the Dai savings rate, and 3) occasionally changing other system-level parameters. The following are the main system-level parameters (see page 15 of (multicollateraldai, )): (1) the debt ceiling for each token type, i.e., the maximum amount of total debt that can be created using a given collateral type, (2) the stability fee, (3) the liquidation ratio, (4) the liquidation penalty, (5) the duration of collateral auctions, (6) the duration of auction bids, and (7) the minimum auction step size (to prevent bids with a minor increment).
@@
Out-of-Scope Concepts. The following DSS concepts are beyond the scope of this paper: (1) the Dai Savings Rate mechanism, which allows users to deposit Dai to collect excess stablecoins created from the stability fee, (2) debt auctions, which occur when a collateral auction fails to raise enough Dai to cover the CDP’s debt, (3) surplus auctions, which occur when the amount of Dai left in vaults due to auctions and stability fees exceeds some threshold, and (4) emergency shutdowns, which can be triggered after a catastrophic event (e.g., system failure) and during which the total remaining collateral is distributed to Dai holders.
2.3. Related Work
The DSS can be seen as an extension of red-black coins (salehi2021red, ), which feature two token types of which one has reduced price volatility (e.g., Dai tokens vs. CDP debt). In (salehi2021red, ), single-asset price simulations are used to estimate the price of the volatile tokens – the main difference with the analysis in Section 5 is that we estimate failure risks, that we use multiple correlated assets, and that we use a heavy-tailed distribution to better represent token price fluctuations. Price simulations of DSS collateral can also be found in (bhat2021simulating, ), and (klagesmundt2022while, ) presents a stochastic model for a single-asset stablecoin. In (grobys2021stability, ), the relationship between stablecoin and Bitcoin price fluctuations is investigated. In (ma2023stablecoin, ), fiat-backed stablecoins are found to have a trade-off between price stability and risk of system failure.
For secure cross-chain transfers, cross-chain deals (herlihy2021cross, ) provide a modeling framework to describe sequences of cross-chain token transfers. In cross-chain deals, the cost of a participant abandoning the protocol before completion (sore-loser attacks (xue2021hedging, )) can be made explicit – this form of misbehavior can be disincentivized through a premium (engel2021failure, ) paid by the parties that abandon the protocol before termination. Cosmos (Polkadot20, ) and Polkadot (cosmos20, ) are blockchain platforms that were specifically developed with blockchain interoperability in mind, but which do not trivially interoperate with existing blockchains outside their ecosystems. However, a version of CroCoDai could be implemented as a native stablecoin on these platforms with collateral on multiple subchains. Alternatively, collateral on multiple chains can be integrated into CroCoDai through trusted relays or zero-knowledge proofs as discussed in Section 3. Finally, Xclaim (zamyatin2019xclaim, ) claims to provide secure cross-chain transfers, but has limitations when transferring to/from smaller blockchain platforms that lack sufficient liquidity because the custodian must submit additional collateral.
3. CroCoDai: A Cross-Chain Stablecoin
Figure 2. Overview of the existing single-chain DSS.
In this section, we present CroCoDai, which extends the functionality of the single-chain DSS (depicted in Figure 2) to a multi-chain setting (depicted in Figure 3). Our main design goal is a stablecoin that is backed by collateral on multiple blockchains, and whose tokens can be moved securely and efficiently between chains. In the following, we will refer to blockchains that store collateral and stablecoin tokens as coin chains (as per (herlihy2021cross, )). The communication between the coin chains is managed by a relay chain, which may be an existing (coin) chain. The main difference between the single-chain DSS and CroCoDai is the inclusion of relay nodes – depicted with double circles in Figure 3. In CroCoDai, all vaults are local to the coin chains, and cross-chain price oracles are supported to complement the oracles on smaller coin chains. This allows CroCoDai to have lower costs for local operations such as creation and transfer of stablecoins on smaller blockchain platforms (we will discuss this in more detail in Section 6.)
In the following, we first discuss in more detail which DSS operations remain local and which require interaction with the relay chain. We then discuss the main components of the coin and relay chains. Next, we discuss two design variants for the relay chain: the first relies on a quorum of trusted nodes who perform most operations off-chain, and the second uses an existing blockchain. We conclude by listing the security requirements of a cross-chain stablecoin, which will be the subject of the Sections 4 and 5.
3.1. Overview
Local Operations. CroCoDai supports the following local operations. CDP Management: Each coin chain has its own vaults that allow users to create CDPs using tokens on that blockchain. As such, stablecoins can be created locally on minor chains, i.e., without the need for transactions on a more expensive chain, as long as the debt limit on the minor chain is not violated. Since CDPs are local, liquidation auctions and other mechanisms such as savings rates are also processed locally. Local Stablecoin Transfers: Each coin chain keeps track of the balance of stablecoins held by its accounts, so the stablecoins can be exchanged locally between accounts. Local Price Information: Finally, local oracles may provide the vaults with price information.
Cross-Chain Operations. The following operations require relay chain involvement. Cross-chain Transactions: In a cross-chain transaction, stablecoins are burned on one coin chain and an equivalent number is minted on another coin chain, as we discuss in more detail below. Ward Management: Governance layer approval is needed to add/remove addresses that are authorized to call security-critical functions such as mint or burn on another contract – we call such addresses wards as per DSS terminology. Periodically, the governance layer may grant or revoke ward status to/from smart contracts for a variety of reasons – e.g., (1) to add support for an emerging blockchain platform, (2) because a critical bug has been found in an existing smart contract, (3) because a chain has undergone a hard fork that interferes with the functioning of the contract, or (4) to implement a more efficient version of a component. System Parameter Updates: Governance layer approval is also needed to update the system parameters discussed in Section 2.2, e.g., a chain’s debt ceiling or a token’s overcollateralization rate on a chain. Global Price Information: The final function is to provide an additional source of token price information to vaults through relay chain oracles – this particularly benefits smaller blockchains that lack a mature DeFi ecosystem.
The exact mechanism for cross-chain transfers is as follows: a user who holds stablecoins on a coin chain first submits a cross-chain transfer request to a relay node and transfers the tokens for escrow. The request is added to a hashmap where it awaits commitment. The nodes on the relay chain monitor the coin chains, and once they detect a cross-chain transfer request, they sign a commit message. When the coin chain receives the commit message, it commits the transaction and burn the tokens held in escrow. A commit message is also sent to the relay node on the target blockchain, which mints an equivalent number of stablecoins. If the transaction is aborted for any reason, then the stablecoins in escrow are returned to the sender.
Figure 3. Overviews of CroCoDai, a multi-chain stablecoin system based on the DSS. The relays (double circle) are not shared with the single-chain DSS.
3.2. Core Components
Coin Chain Components. Each coin chain has the following components. Coin: implements a contract that is similar to Dai in the DSS, and processes transfers of stablecoins between users on the coin chain. There is only one coin contract per chain. Vault: implements contracts that are similar to Vat, GemJoin and DaiJoin in the DSS, which allow users to create CDPs and withdraw/deposit collateral and stablecoins. There can be multiple vaults per chain. Auction: implements auction contracts that liquidate the collateral of underwater CDPs, with any appropriate auction logic – e.g., Cat or Dog in the DSS. There can be multiple auction contracts per chain. Bottom-Level Oracle: implements a price feed using a contract similar to Spot in the DSS, combined with a Medianizer. Bottom-Level Relay: validates messages from the relay chain. It processes (1) the addition of wards to CroCoDai contracts, (2) incoming cross-chain transfers, (3) commitments of outgoing cross-chain transfers, and (4) updates of system parameters.
Relay Chain Components. The relay chain has the following components. Governance: tracks the ownership of governance tokens, and allows users to transfer them. As in the DSS, users who hold governance tokens can vote on system-level actions such as adding wards and updating system-level parameters. Our security assumption regarding governance token holders is that a majority is honest and online at any point during the voting window for system-level decisions. Top-Level Oracle: The relay chain may implement system-level price feeds to complement the price feeds on the coin chains. Top-Level Relay: To process cross-chain transfers, the relay chain must be able to process information from the coin chains, particularly the transfer requests – we discuss this in more detail below.
3.3. Relay Chain Designs
In this section, we discuss two possible design variants of the relay chain. The first relies on a quorum of trusted nodes with known public keys – each trusted node runs an off-chain (full or light) blockchain client for each of the underlying chains to check the validity of cross-chain transfers, and the nodes maintain a distributed ledger for governance decisions. The second uses smart contracts on an existing chain to implement the consensus logic of the underlying chains (potentially using zero-knowledge cryptography) to verify cross-chain operations on-chain.
Off-Chain Blockchain Clients. In this approach, the stablecoin system is maintained by $n$ relay nodes with known public keys. The relay nodes use their (light) clients to detect and validate cross-chain transactions, and generate signatures using their private/public key pair to commit cross-chain transfers and process system-level actions approved by governance token holders. In this setting, the only functionality of the bottom-chain relay contracts is to verify digital signatures. We discuss three different subvariants of this approach in Section 6. As is common in Byzantine fault tolerant systems (castro1999practical, ; lamport1982byzantine, ), we assume that at most $f$ nodes are faulty or malicious such that $n\geq 3f+1$. We note that it is vital that the relay nodes are controlled by independent entities: e.g., in the attack against the Ronin bridge, 5 out of 9 validators were successfully attacked (roninhack, ), of which 4 belonged to the same entity (roninhack2, ).
On-Chain Consensus Logic. In this approach, the consensus logic of the coin chains is implemented in the top-level relay contracts, and the consensus logic of the relay chain is implemented in the bottom-level relay contracts. Essentially, the functionality of a light client – i.e., validating block headers and proving the state of a contract using the header – is implemented inside the smart contracts. This is the approach taken in, e.g., Xclaim (zamyatin2019xclaim, ). For example, in Bitcoin’s Nakamoto Consensus (btc_origin, ) this means that the smart contract computes header hashes and compares their integer value to the proof-of-work difficulty, [^4] This may be computationally efficient if blockchains share built-in hash functions, otherwise this can be expensive. whereas in Proof-of-Stake Ethereum the smart contract verifies that the total stake of the nodes who have approved a block is sufficient. In this setting, the main role of the relay nodes is to generate and submit validity proofs. Techniques to increase the efficiency of light clients have been proposed in the literature for both proof-of-work (e.g., Bitcoin (daveas2020gas, )) and proof-of-stake blockchains (e.g., Algorand (micali2021compact, ) and Mina (bonneau2020mina, )) – in some cases using zero-knowledge proofs that may incorporate the blockchain’s consensus logic (xie2022zkbridge, ). Zero-knowledge proofs that are verifiable on multiple chains, e.g., ZKM’s entangled rollups (zkm2024entangled, ), can also be used to bridge the chains that implement collateral to a layer-2 network that implements the vaults. Finally, the data on the different coin chains can be integrated using TrustBoost (wang2022trustboost, ).
Although the second design variant avoids a trust assumption on a quorum of relay nodes, current iterations of such designs have a large computational overhead, typically requiring several kilobytes of data sent to relay contracts and more than 10 seconds to generate signatures (micali2021compact, ; xie2022zkbridge, ). [^5] Improving the performance of such zero-knowledge schemes, e.g., using folding schemes (kothapalli2022nova, ), is an active research area. As such, we focus on the first variant in Section 6 and leave the second as future work.
3.4. System Requirements